# Easy conditioning far beyond Gaussian

### v3

@@ 4 Implementation of the method
Our method’s versatility stems from its ability to choose any dependence structure among families of copulas stable by conditioning, which includes mixtures, and to model the joint distribution with arbitrary marginals. The primary challenge in our procedure lies in fitting a copula in the uniform space. This involves balancing the complexity of the dependence structure against the precision with which we can estimate the copula parameters. For this paper, we have opted to use a Gaussian mixture copula with a limited number of components. However, alternative dependence structures such as Student’s $t$-mixture copulas or Skew-Normal mixture copulas represent exciting avenues for future research.
@@ 4.1 Gaussian Mixture Copula Model
The Gaussian mixture copula [27] is the copula implicit in a Gaussian mixture model. It allows to capture multi-modality in the dependence structure. The density of a Gaussian mixture model (GMM) with $K$ components is given by:
@@
Similarly, we can define the Student $t$-mixture copula [17] or the unified Skew-Normal mixture copula as others implicit copulas.
@@ 4.1.1 Fitting of a GMCM
Estimating parameters of a GMCM is a challenging task for several reasons as it has been discussed in [26].
The first challenge is that copulas are invariant under marginal increasing transformations. Specifically, if $\mathbf{X}=(X_{1},\ldots,X_{d})^{t}$ is a random vector with continuous marginal distributions and copula $C$, and $T_{1},\ldots,T_{d}$ are strictly increasing functions. Then $(T_{1}(X_{1}),\ldots,T_{d}(X_{d}))^{t}$ also has copula $C$.
In the case of a GMCM, this property implies a lack of identifiability of the parameters, which makes the fitting of the parameters more difficult. This can be seen through the following proposition [26]:
Proposition 4.1.
@@
This lack of identifiability results in the likelihood of a GMCM having infinitely many maximizers. To address this issue, Bilgrau [6] proposed standardizing the first component to have zero mean and unit variance along each dimension, which resolves the identifiability issue in most cases.
Another challenge in fitting GMCM is the high computational cost associated with functional evaluations, as the quantile function of a Gaussian mixture model isn’t known in closed form. To address this, we can use numerical approximation schemes leveraging the monotonicity of the cdf, although they tend to be computationally expensive. Alternatively, methods based on the bisection algorithm, such as the one developed by Chandrupatla [10], can be employed to find the inverse of a cdf. Since the derivatives of the Gaussian mixture model’s cdf are known analytically, a Newton-type algorithm can also be effectively used in this context.
In the literature, two primary methods have been proposed for estimating the parameters of a GMCM via direct gradient-based optimization. The classical approach involves using finite difference methods (denoted as FD in the rest of this paper) to estimate the derivatives. This approach approximates the derivatives numerically, which can be computationally expensive and less precise.
@@
The methods FD and PEM are implemented in the R package GMCM [6]. Meanwhile, the AD method has been developed in Python by Tewari.
@@ 4.1.2 Comparison of the copula fitting methods
To compare the different optimization methods described above (AD, FD and PEM), we simulated data following a GMCM with known parameters and attempted to recover its values using different fitting procedures. Because GMCM parameters are not identifiable, we cannot directly compare the recovered parameter values. Instead, we need to compare metrics such as the log-likelihood or distances to the true distribution. We used the Cramér distance to quantify the differences between empirical multivariate distributions. We compared samples of the true underlying distributions to samples of the distribution with the fitted parameters and repeated the experiments several times to take into account the uncertainty in the sampling of a GMCM. In the results, we indicated the average Cramér distance and average log-likelihood as well as the standard deviations.
@@ 5 Numerical experiments
We split our data set randomly into a training set containing $80\%$ of the observations and a test set containing the remaining $20\%$ of the observations. We use the training set to learn the joint multivariate distribution and we sample $1000$ points from the estimated conditional distribution at every test points.
We then evaluate the performance of our approach comparing the conditional samples to the observed values using scoring rules. In the univariate case, we use the Continuous Ranked Probability Score (CRPS) and the logarithmic score. For multivariate outcomes, we employ the Energy score and the Variogram score [16], with order $r=0.5$. In particular, we will examine the mean value of the various scores across all test points, averaged over multiple train/test splits.
@@
Given the complexity of fitting a Gaussian Mixture Copula Model (GMCM) in the uniform space, an alternative approach for conditional sampling involves fitting a Gaussian Mixture Model (GMM) in a transformed space using the EM algorithm, which boasts theoretical convergence guarantees. Specifically, we can transform each marginal distribution to an approximate standard normal distribution. Following this, we can perform conditioning within this space and subsequently transform the samples back to the original space. We will refer to this method as Transformed Gaussian Mixture Model (TGMM).
@@ 5.1.1 Simulation scenarios
We consider two simulation scenarios where the response is univariate:
@@
    $X_{1}|X_{2}=x_{2}\sim\tilde{\alpha}_{1}(x_{2})\mathcal{N}(4+(x_{2}-2),1)+\tilde{\alpha}_{2}(x_{2})\mathcal{N}(\frac{(x_{2}-1)}{2}-2,0.75)$
@@
    In this test case, the data follows a Gaussian mixture dependence structure, and we know the specific number of components. As a result, we anticipate that our method GMCM, will perform very well.
@@
    Let’s now consider another example with a univariate response, where the data has been simulated from a $2$-dimensional meta-Gaussian mixture model. This means that the copula corresponds to that of a Gaussian mixture model, while the marginal distributions can have various forms.
    We consider a Gaussian mixture copula model with the same parameters as in the previous example. The marginals are chosen to be standard normal distributed. An illustration of this multivariate distribution can be found in figure 7.
@@ 5.1.2 Results
Table 3: Comparison of various methods based on the mean of scoring rules over the test set for the various simulation scenarios.
@@
GC 3.141 3.055 0.875 2.056
GMCM 1.277 1.860 0.436 1.172
TGMM 1.279 1.889 0.440 1.177
CKDE 1.367 2.090 0.530 1.479
The results indicate that our method using Gaussian Mixture Copula Model outperforms the competitors for conditional sampling on simple examples.
Notice that another advantage of our method is that it provides estimates of the cdf $\hat{F}_{\mathbf{X_{1}}|\mathbf{X_{2}}=\mathbf{x_{2}}}(\mathbf{x_{1}})$ which are both smooth in $\mathbf{x_{2}}$ but also in $\mathbf{x_{1}}$. However, we need to fix the number of mixtures components beforehand, which would require some prior knowledge about the data. One potential solution is to select a large number of mixture components to enhance the modeling capacity of our model, thereby capturing more complex dependence structures. Nonetheless, this approach introduces greater challenges in parameter estimation. Consequently, the advantage gained from a more comprehensive modeling of the joint distribution is offset by the reduced accuracy in parameter estimation.
@@ 5.2 Real world examples
We now turn to real datasets with multivariate outputs and use multivariate scoring rules for evaluation.
@@ 5.2.1 Wine dataset
This dataset already presented in the introduction contains the amount of $13$ constituents found in each of the three types of wines and and consists of $178$ observations. We will set the number of components in our GMCM to 3. For our analysis, we will work with a $5$-dimensional problem where we try to predict the $3$ first variables from joint observations of the $5$ first variables. A graphical representation of this dataset can be found in figure 9.
The results, in table 4, show that the GMCM performs similarly than GC and TGMM. However, the CKDE benchmark seems to outperform all the other methods on this example. This could be attributed to the modeling of marginal distributions, for which we assume a Gaussian mixture form in our methodology, and seems not to be adequate in this example.
Table 4: Comparison of various methods based on the mean of scoring rules over the test set for the Wine dataset.
@@
GC 1.254 0.470
GMCM 1.221 0.489
TGMM 1.308 0.491
CKDE 1.207 0.466
@@ 5.2.2 Breast Cancer data Wisconsin
The data set provides data for $569$ patients on $30$ features of the cell nuclei obtained from a digitized image of a fine needle aspirate (FNA) of a breast mass. For each patient the cancer was diagnosed as malignant or benign. This data-set has been studied for classification or clustering tasks.
In this section, we will work again with a $5$ dimensional problem by predicting the $3$ first variables from joint observations of the $5$ first variables. A graphical representation of the data-set can be found in figure 11.
Given that the dataset appears to contain two clusters corresponding to malignant and benign diagnoses, we will set the number of components in our GMCM to 2. The results in table 5 demonstrate that our method outperforms the other methods for conditional sampling on this example.
Table 5: Comparison of various methods based on the mean of scoring rules over the test set for the breast cancer data Wisconsin.
@@
GC 1.896 0.848
GMCM 1.845 0.823
TGMM 1.854 0.826
CKDE 2.007 1.979
@@ 6 Conclusion and perspectives
In this work, we have proposed an interpretable method allowing easily estimating and sampling from conditional distributions based on estimated joint distributions belonging to a broad class way beyond the Gaussian case. We introduced the notion of trans-dimensional families of multivariate distributions that are stable by marginalization and conditioning. This class encompasses well-known families of multivariate distributions such as multivariate Gaussian distributions, Student t-distributions, and skew normal distributions. We demonstrated that this class extends to mixtures and transformations, providing a wide range of multivariate distributions that remain stable by conditioning.
Leveraging this stability property, we focused on distributions with tractable formulas for their conditional distributions. We developed an algorithm to estimate or sample from conditional distributions by modeling the joint distribution with a meta distribution. This approach uses copulas that are stable by conditioning with arbitrary marginals. While our method is highly versatile, for practical reasons, we concentrated in applications on the Gaussian copula mixture model. In turn, we empirically compared different approaches to the associated copula parameter fitting problem.
@@
Numerical results indicate that our approach can accurately estimate conditional distributions in simple synthetic cases. On examples with real datasets, the performance of our method is more variable and sometimes lower than the benchmarks. However, in cases where the data exhibits a multi-modal nature, leveraging multiple components within the Gaussian mixture copulas framework can offer significant benefits. Furthermore our method is interpretable and preserves privacy, which is not the case of the other benchmarks. In particular, comparing different classes of imputation (as well as synthetic generation) methods both in terms of scoring approaches such as used here and of privacy indicators would be of great interest for future work.
@@
Given the promising results of our method in the considered continuous framework, an important next step is to extend this methodology to handle mixed variables, which include both continuous and discrete components. Such an extension would be particularly relevant in medical applications where health outcomes of interest are often measured as categorical or dichotomous states. A possible route would be to to assume the discrete variables are discretizations of latent continuous variables as is common practice in latent class modelling frameworks.
@@
Future research directions also include extending our method by implementing the algorithm for other dependence structures, such as mixtures of Student t-distributions and mixtures of skew normal distributions. Additionally, we aim to extend our work to accommodate missing values in the observed sample.
Finally, a deeper exploration of the theoretical properties of the proposed method could yield new insights and potentially pave the way for novel applications across various fields.
@@
The code used to estimate and sample from the conditional distributions is available in Python. The datasets are publicly available online.
Wine data: https://archive.ics.uci.edu/ml/machine-learning-databases/wine/wine.data
Breast Cancer data: https://archive.ics.uci.edu/dataset/17/breast+cancer+wisconsin+diagnostic
Disclosure Statement
The authors report that there are no competing interests to declare.
@@ Appendix A Multivariate Student $t$-distribution
The multivariate Student $t$-distribution is stable by marginalization and conditioning as states the two following propositions [11]. In this appendix, we will use fixed, ordered sets of indices, with the first set containing the variables we wish to condition on the variables in the second set. These results can be readily extended to any arbitrary set of indices.
Proposition A.2.

### v4

@@ 4 Implementation of the method
Our method’s versatility stems from its ability to choose any dependence structure among families of copulas stable by conditioning, which includes mixtures, and to model the joint distribution with arbitrary marginals. The primary challenge in our procedure lies in fitting a copula in the uniform space. This involves balancing the complexity of the dependence structure against the precision with which we can estimate the copula parameters. For this paper, we have opted to use a Gaussian mixture copula with a limited number of components. Usng further dependence structures such as Student’s $t$-mixture copulas or Skew-Normal mixture copulas represent exciting avenues for future research.
@@ 4.1 Gaussian Mixture Copula Model
The Gaussian mixture copula [tewari2011parametric] is the copula implicit in a Gaussian mixture model. It allows to capture multi-modality in the dependence structure. The density of a Gaussian mixture model (GMM) with $K$ components is given by:
@@
Similarly, we can define the Student $t$-mixture copula [massing2021clustering] or the unified Skew-Normal mixture copula as others implicit copulas.
@@ 4.1.1 Fitting of a GMCM
Estimating parameters of a GMCM is a challenging task for several reasons as it has been discussed in [tewari2023estimation].
The first challenge is that copula invariance under marginal increasing transformations causes a lack of identifiability of the parameters.
This can be seen through the following proposition [tewari2023estimation]:
Proposition 1.
@@
This lack of identifiability results in the likelihood of a GMCM having infinitely many maximizers. To address this issue, Bilgrau [bilgrau2016gmcm] proposed standardizing the first component to have zero mean and unit variance along each dimension.
Another challenge in fitting GMCM is the high computational cost associated with functional evaluations, as the quantile function of a Gaussian mixture model isn’t known in closed form. To address this, we can use numerical approximation schemes leveraging the monotonicity of the cdf, although they tend to be computationally expensive. Alternatively, methods based on the bisection algorithm, such as the one developed by Chandrupatla [chandrupatla1997new], can be employed to find the inverse of a cdf. Since the derivatives of the Gaussian mixture model’s cdf are known analytically, a Newton-type algorithm can also be effectively used in this context.
In the literature, two primary methods have been proposed for estimating the parameters of a GMCM via direct gradient-based optimization. The classical approach involves using finite difference methods (denoted as FD in the rest of this paper) to estimate the derivatives. Approximating derivatives numerically can cause imprecisions and heavy computational costs.
@@
The methods FD and PEM are implemented in the R package GMCM [bilgrau2016gmcm].Meanwhile, the AD method has been developed in Python by Tewari in a Github page related to this paper [tewari2023estimation].
@@ 4.1.2 Comparison of the copula fitting methods
To compare the different optimization methods described above (AD, FD and PEM), we simulate data following a GMCM with known parameters and attempted to recover its values using different fitting procedures. Because GMCM parameters are not identifiable, we cannot directly compare the recovered parameter values. Instead, we chose to compare metrics such as the log-likelihood or distances to the true distribution. We used the Cramér distance to quantify the differences between empirical multivariate distributions. We compared samples of the true underlying distributions to samples of the distribution with the fitted parameters and repeated the experiments several times to take into account the uncertainty in the sampling of a GMCM. In the results, we indicated the average Cramér distance and average log-likelihood as well as the standard deviations.
@@ 5 Numerical experiments
We split the considered data sets randomly into a training set containing $80\%$ of the observations and a test set containing the remaining $20\%$ of the observations. We use the training set to learn the joint multivariate distribution and we sample $1000$ points from the estimated conditional distribution at every test points.
We then evaluate the performance of our approach comparing the conditional samples to the observed values using scoring rules. In the univariate case, we use the Continuous Ranked Probability Score (CRPS) and the logarithmic score. For multivariate outcomes, we employ the Energy score and the Variogram score [jordan2017evaluating], with order $r=0.5$. In particular, we will examine the mean value of the various scores across all test points, averaged over multiple train/test splits.
@@
Given the complexity of fitting a GMCM in the uniform space, an alternative approach for conditional sampling involves fitting a GMM in a transformed space using the EM algorithm, which boasts theoretical convergence guarantees. Specifically, we can transform each marginal distribution to an approximate standard normal distribution. Following this, we can perform conditioning within this space and subsequently transform the samples back to the original space. We will refer to this method as Transformed Gaussian Mixture Model (TGMM).
@@ 5.1.1 Simulation scenarios
We consider two simulation scenarios focused on sampling from a univariate conditional distribution:
@@
    $X_{1}|X_{2}=x_{2}\sim\tilde{\alpha}_{1}(x_{2})\mathcal{N}(4+(x_{2}-2),1)+\tilde{\alpha}_{2}(x_{2})\mathcal{N}\left(\frac{(x_{2}-1)}{2}-2,0.75\right)$
@@
    In this test case, the data follows a Gaussian mixture dependence structure, and we know the specific number of components.
@@
    We now consider another example with a univariate response, where the data has been simulated from a $2$-dimensional meta-Gaussian mixture model. Here the marginal distributions are chosen standard normal, and the dependency structure is driven by a Gaussian mixture copula model with the parameters $\alpha_{1}=\alpha_{2}=0.5$ and
    $\mu_{1}=\begin{pmatrix}2\\ 2\end{pmatrix},~\mu_{2}=\begin{pmatrix}-2\\ -2\end{pmatrix}$ and $\Sigma_{1}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix},~\Sigma_{2}=\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$.
    An illustration of this multivariate distribution can be found in figure 7.
@@ 5.1.2 Results
Table 3: Comparison of various methods based on average scores over the test set for the various simulation scenarios.
@@
GC 1.207 1.878 0.328 0.922
GMCM 1.193 1.832 0.300 0.700
TGMM 1.214 1.864 0.318 0.870
CKDE 1.304 2.066 0.431 1.333
The results indicate that our method using GMCM and the TGMM outperform the other competitors for conditional sampling on simple examples.
In the next section, we turn to real datasets with multivariate outputs.
@@ 5.2.1 Wine dataset
This dataset already presented in the introduction contains the amount of $13$ constituents found in each of the three types of wines and consists of $178$ observations. We set the number of components in our GMCM to 3. For our analysis, we work with a $5$-dimensional problem where we try to predict the $3$ first variables given the $2$ remaining ones, based on training data consisting of from joint observations of the $5$ considered variables. A graphical representation of this dataset can be found in figure 9.
The results, in table 4, show that the GMCM performs better than GC and TGMM in terms of Energy Score (scoring rule in multivariate settings that generalizes CRPS). However, the CKDE benchmark seems to outperform all the other methods on this example.
Table 4: Comparison of various methods based on average scores over the test set for the Wine dataset.
@@
GC 1.263 0.504
GMCM 1.216 0.491
TGMM 1.290 0.493
CKDE 1.211 0.468
@@ 5.2.2 Breast Cancer data Wisconsin
The Breast Cancer Wisconsin dataset sourced from the UCI Machine Learning Repository, utilizes digitized images from fine needle aspirates (FNA) of breast masses. It comprises ten features derived from $569$ patients. For each feature, the dataset provides the mean value, extreme value (average of the three largest values), and standard error, resulting in $30$ continuous variables. The dataset includes two types of diagnoses: benign ($352$ patients) and malignant ($212$ patients), suggesting the presence of two distinct clusters. For this experiment, we focus on a subset of four variables: perimeter standard error (PSE), extreme smoothness (ES), extreme concavity (EC), and extreme concave points (ECP).
We aim to use probabilistic methods to predict the values of EC and ECP based on observations of PSE and ES.
A graphical representation of the data-set can be found in figure 11.
Given that the dataset appears to contain two clusters corresponding to malignant and benign diagnoses, we will set the number of components in our GMCM to 2. The results in table 5 demonstrate that our method and the TGMM perform competitively on this example.
Table 5: Comparison of various methods based on average scorings over the test set for the breast cancer data Wisconsin.
@@
GC 0.725 0.327
GMCM 0.671 0.285
TGMM 0.669 0.281
CKDE 1.012 0.523
5.3 Probabilistic imputation
One key advantage of the proposed approach is its flexibility: once the joint distribution is learned, we can analytically compute the conditional distribution for any combination of observed and missing variables This property is particularly valuable for imputation tasks, where different variables may be missing for different observations. In contrast, alternative distributional regression methods typically require fitting a separate model for each pattern of missing and observed variables.
Starting with a complete dataset, we randomly introduce missingness in a given proportion of entries (here, $10\%$). Observations with no missing values form the training set, while those with one or more missing values comprise the test set. We fit a Gaussian Mixture Copula Model (GMCM) to the training data. For each test observation, we sample from the conditional distribution of its missing variables, conditioned on its observed values. This produces multiple imputed datasets.
To evaluate imputation quality, we employ scoring rules. At first glance, applying scoring rules appears challenging: typically, one calculates an average score across all test predictions, but in our scenario, predictions may have different dimensions, complicating fair comparison.
One approach to circumvent this issue is to "complete" all predictions to joint distributions by plugging in the true observed values. However, it is worth questioning if evaluating these completed joint samples aligns with our actual imputation goal. Notably, for the Energy Score, it turns out that evaluating either the joint samples or the conditional samples of varying dimension is equivalent. In particular, for univariate predictions, this process reduces to calculating the CRPS.
We compared our approach with MICE [van2011mice], a widely used multiple-imputation procedure using chained equations. Our method was applied to a subset of four variables from the Breast Cancer Wisconsin dataset, see section 5.2.2, known for its multimodal structure, as well as a simulated dataset generated from a GMCM. For further details, we refer to Appendix C.3.
The results in Tables 7 and 6 indicate that GMCM slightly outperforms MICE in terms of Energy Score while requiring significantly less computational time. In experiments with a small number of features, our method performs comparably to MICE yet consistently shows a clear computational advantage. However, as the dimensionality increases, MICE tends to perform better. These findings suggest that our algorithm is well-suited for imputation tasks in datasets with few features and a multimodal structure, providing competitive results to MICE at a reduced computational cost.
Table 6: Imputation results for the Wisconsin data set
Metric GMCM MICE
Energy Score 0.252 0.286
Computational Time (s) 102.1 180.7
Table 7: Imputation results for simulated data
Metric GMCM MICE
Energy Score 0.097 0.102
Computational Time (s) 158.3 230.03
@@ 6 Conclusion and perspectives
In this work, we explored families of multivariate distributions that allow for analytical computation of conditional distributions. We provided examples beyond the Gaussian case, such as the multivariate Student $t$-distribution or the multivariate skew normal distribution.
Additionally, we introduced the concept of trans-dimensional families of multivariate distributions that are stable by conditioning. We demonstrated that this class extends to finite mixtures and transformations, showing that a broader range of multivariate distributions can benefit from analytical conditioning.
Leveraging these results, we developed a versatile approach for estimating conditional distributions from data. In practice, we model joint distributions using meta distributions with copulas that are stable by conditioning. In these settings, conditioning in original space is performed via a latent space where conditioning is more straightforward.
While our method allows for a high flexibility when considering broader families enjoying stability by conditioning, we focused in applications on the Gaussian Mixture Copula Model (GMCM) for practical purposes. In turn, we empirically compared different approaches to tackle the associated copula parameter fitting problem.
@@
Numerical results indicate that our approach can accurately estimate conditional distributions in simple synthetic cases. Experiences on real datasets shows that the performance of our method is more variable and sometimes lower than the benchmarks. However, in cases where the data exhibits a multi-modal nature, leveraging multiple components within the GMCM framework can offer significant benefits. Moreover, our method offers several advantages over competitors: it is interpretable, provides an analytical form for the conditional distribution, and enables amortized inference for any combination of missing variables, which motivates its application to imputation tasks.
@@
Given the promising results of our method in the considered continuous framework, an important next step is to extend this methodology to handle mixed variables, which include both continuous and discrete components. Such an extension would be particularly relevant in medical applications where health outcomes of interest are often measured as categorical or dichotomous states. A possible route would be to assume the discrete variables are discretizations of latent continuous variables as is common practice in latent class modeling frameworks. One limitation of our work is the challenge of fitting copulas, especially as dimensionality increases. Specifically, for the GMCM, the number of parameters scales quadratically with dimension, making it feasible only for relatively low-dimensional problems. Moreover, we currently need to manually preset the number of mixture components in the GMCM, whereas an automated method for determining this would be beneficial.
In higher-dimensional settings, other generative methods for sampling from a conditional density use tensor-train decomposition to approximate the joint density and sample from the conditional distribution by using the inverse Rosenblatt transform [cui2023scalable]. This approach has been shown to have strong theoretical guarantees, however it requires a predetermined ordering of variables, which can limit its usefulness in tasks like imputation.
@@
Future research directions include extending our method by implementing the algorithm for other dependence structures, such as mixtures of Student t-distributions and mixtures of skew normal distributions. Additionally, we aim to extend our work to accommodate missing values in the observed sample.
@@
The code used in this paper has been developed in Python and is publicly available on GitHub. The datasets are publicly available online.
@@ Appendix A Multivariate Student $t$-distribution
The multivariate Student $t$-distribution is stable by marginalization and conditioning as states the two following propositions ([ding2016conditional]). In this appendix, we will use fixed, ordered sets of indices, with the first set containing the variables we wish to condition on the variables in the second set. These results can be readily extended to any arbitrary set of indices.
Proposition 2.